OpenStudy (anonymous):

Kit folds a bandana diagonally before tying it around her head. The side length of the bandana is 16 in. About how long is the diagonal

OpenStudy (anonymous):

 a^2 + b^2 = C^2 ----> 16^2 + 16^2 = x^2.